OpenStudy (anonymous):

find the slope and y- intercept of the line whose equation is 4x+y=9

OpenStudy (anonymous):

given eq is 4x+ y=9 it can also be written as y= -4x+ 9 comparing it with y = mx+c we find slope m= -4 and y-intercept c= 9 @ken_boy05

OpenStudy (anonymous):

Thanks a bunch ...im learning a lot

OpenStudy (anonymous):

@ken_boy05 hey ken don't forget to award the medal...by clicking on the right side button along side my name.

OpenStudy (anonymous):

sure ...are u sure the y intercept is 9?

OpenStudy (anonymous):

yes 100%

OpenStudy (anonymous):

ok i was trying to write it in ordered pair and it was saying wrong but i figured it out (0,9)

OpenStudy (anonymous):

another way to find y intercept just plug x= 0 in the given eq of line u will get y= 9 which is y-intercept

OpenStudy (anonymous):

i have another one will close this and pput it up in a new window see if u can get another medal

OpenStudy (anonymous):

sure